OpenStudy (anonymous):

Can someone show the solution to this DE? x*dy/dx-4y = (e^x)*(x^4)

OpenStudy (anonymous):

divide by x and you will have an equation in the form \[\frac{ dy }{ dx } - \frac{4}{x}y = (e^{x})(x^{4})\frac{1}{x}\]

OpenStudy (anonymous):

do you know how what an integrating factor is?

OpenStudy (anonymous):

yeah, i tried it, i got all messed up. can you show me?

OpenStudy (anonymous):

what did you have as the integrating factor?

OpenStudy (anonymous):

4/x and x^3 * e^x

OpenStudy (anonymous):

so the integrating factor is \[p = \exp(\int\limits_{}^{} g(x) dx)\]

OpenStudy (anonymous):

where in your case \[g(x) = \frac{-4}{x}\]

OpenStudy (anonymous):

\[\frac{dy}{dx} + g(x)y = h(x)\] \[p = \exp(\int\limits\limits_{}^{} g(x) dx)\] where p is the integrating factor. times everything through by p \[p\frac{dy}{dx} + pg(x)y = ph(x)\] now you can rewrite the left hand side because: \[\frac{d}{dx}(p(x)y) = p\frac{dy}{dx} +pg(x)y\]

OpenStudy (anonymous):

and finally integrating gives \[p(x)y = \int\limits_{}^{}p(x)h(x)dx\]

OpenStudy (anonymous):

divide by p(x) to give y in terms of an integral. Integrate and you have solved for y